# Trust-Aware Assistance-Seeking in Human-Supervised Autonomy

### v2

@@ 1. Introduction
Autonomous systems are becoming increasingly prevalent and are being deployed and used in various sectors such as healthcare, agriculture, military, and transportation. Despite the increasing capabilities of autonomous systems, these systems still require human supervision and assistance when working in complex and uncertain environments(Selma2017, ; akash2020human, ; peterssupervisor2015, ; gupta2024structural, ; Boguslavskii2025, ). However, the increased capabilities of the robots are leading to novel landscapes of human-robot interaction (HRI) wherein the robots are deployed in roles requiring complex social behavior with human partners.
@@
Human trust in the robotic system is a key factor to be considered for effective collaboration between humans and robots in complex and uncertain environments. While trust has been traditionally studied in the context of collaborative human teams (mayer1995integrative, ), its influence in heterogeneous human-robot teaming is gaining prominence with the increased capabilities and deployment of robotic systems. Lee and See (lee2004trust, ) define trust in an autonomous agent as “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability”. Wagner and Arkin (wagner2011recognizing, ) generalize this definition to define trust as “belief held by the trustor that the trustee will act in a manner that mitigates the trustor’s risk in a situation in which the trustor has put its outcome at risk.” Trust in a robotic system is not only influenced by its physical design but also by the autonomous decisions and actions executed by the robot.
@@
While having low trust is undesirable as it leads to the disuse of a system, having high trust can also be undesirable as it can lead to misuse or abuse of the system (hoff2015trust, ; hancock2011meta, ; parasuraman1997humans, ). Therefore, for an effective design of human-robot teams, the calibration of the human partner’s trust in the robot’s abilities and capabilities is necessary. To this end, it is vital to understand the influence of various robot-related, environment-related, and human-related factors on the dynamic evolution of trust.
@@
The classical approach to studying the impact of robot reliability, transparency, and human workload on human trust treats trust as a static parameter (hoff2015trust, ). However, recent studies have focused on understanding the dynamic evolution of human-automation trust, focusing on factors such as the quality of robot performance, transparency of the robot’s operation, and the human agent’s attitude and characteristics, such as forgetfulness and self-confidence, on the trust dynamics (lee1992trust, ; lee1994trust, ; manzey2012human, ; desai2013impact, ; akash2017dynamic, ; yang2017evaluating, ; Wang2015DynamicRS, ; chen2020trust, ; akash2020human, ; williams2023computational, ). Hancock et al. (hancock2011meta, ) present a comprehensive review of various factors affecting human trust and categorize them into three major groups: robot-related, human-related, and environment-related factors. They discuss that human trust is most influenced by robot-related factors, which include robot reliability, performance, design, and abilities. This is followed by the environment-related factors, which include task complexity, type, and physical environment. Among the three main categories, human-related factors such as demographics, expertise, and personality have the least influence on trust in human-robot interaction.
@@
Deterministic linear models use a linear time-invariant dynamical system, where the states are the current trust level, cumulative trust level, and expectation bias (akash2017dynamic, ; yang2017evaluating, ; Wang2015DynamicRS, ). The input to these linear models includes human experience, which is determined by the performance and reliability of the robot. Variants of these models have been introduced to capture mutual trust between humans and robots (rahman2018mutual, ; Wang2014, ). These models have been used in several contexts, including exploring the impact of information transparency in reconnaissance missions (yang2017evaluating, ), investigating human reliance on driving assistance systems (akash2017dynamic, ), examining subtask allocation in collaborative assembly tasks (rahman2018mutual, ), and addressing scheduling challenges in teleoperation for underwater robotic navigation (Wang2014, ).
@@
Stochastic models of trust often take one of two forms. Stochastic linear models for the evolution of trust (azevedo2021real, ) rely on linear time-invariant dynamical systems with Gaussian measurement noise. In contrast, other stochastic models treat trust as a categorical latent variable and estimate its distribution based on observable factors like robot performance and human actions (OPTIMO, ; liao2024trust, ). An example of this class of models is the Input-Output Hidden Markov Model (IOHMM) with categorical states, which have been previously used to model human trust (akash2020human, ; DBLP:journals/corr/abs-2009-11890, ; williams2023computational, ). These models are particularly suitable for capturing interpretable internal states (e.g., “High Trust” or “Low Trust”). In addition, unlike linear Gaussian systems, which typically rely on richer or continuous signals such as self-reported trust on Likert scales (chen2020trust, ), gaze duration, or attention time (azevedo2021real, ), or the presence of a secondary task, the IOHMM provide flexibility for handling discrete observations such as human decisions (e.g., Rely or Intervene, Yes or No, Agree or Disagree) in both training and prediction. A comprehensive survey of different trust models is presented in (wang2023human, ). To optimize robot decision-making under uncertainty about internal human states, Partially Observable Markov Decision Process (POMDP) framework has been employed with trust as latent state influenced by robot actions, human responses, and contextual factors (chen2020trust, ; zahedi2023trust, ; akash2020human, ; 10.5555/2906831.2906852, ; DBLP:journals/corr/abs-2009-11890, ). It has been used in applications including scheduling in human-supervised robotics to reduce human interventions (chen2020trust, ) and monitoring load (zahedi2023trust, ), as well as generating optimal recommendations and explanations based on inferred human state (akash2020human, ; 10.5555/2906831.2906852, ). In this work, we combine IOHMM-based trust modeling with a POMDP-based decision-making framework.
@@
In addition to a dynamic trust model, real-time estimation of trust is vital for the design of trust-based policies that are robust to modeling uncertainties. Traditionally, human trust is assessed through self-reporting, such as surveys (schaefer2016measuring, ; lee1992trust, ; muir1989operators, ). However, a limitation of this approach is its impracticality for frequent trust survey completion during the experiment. Psycho-physiological measurements(gebru2022review, ) have also been used to infer the trust state; however, the sensing devices required to collect these measurements, e.g., an EEG head cap, limit their applicability. For a more detailed review of trust measurements, refer to (gebru2022review, ). These challenges have prompted researchers to rely purely on context and behavioral data, such as human actions, to estimate trust. This approach has been implemented and tested in several contexts such as reconnaissance mission (akash2020human, ), autonomous driving (azevedo2021real, ; DBLP:journals/corr/abs-2009-11890, ), and supervision(OPTIMO, ). We focus on similar methodology for estimating human trust: subjective measures and behavioral data.
@@
In this paper, in the context of human-supervised object collection tasks, we design and validate trust-aware policies for an autonomous system seeking assistance from a human supervisor. In contrast to existing works, our focus is on a scenario where the robot can seek assistance from the human supervisor, preventing a loss of trust and possibly repairing it. The major contributions of this work are the following. First, we develop a novel experiment to explore how a robot’s requests for assistance influence human trust. Second, we use an Input-Output Hidden Markov Model (IOHMM) (536317, ) to capture the evolution of the hidden human trust state. In our formulation, the inputs are the robot’s action, environmental complexity, and the outcome or experience in the previous trial, characterized by the success or failure of the robot in accomplishing its goals, and the output is the human action. The present research sought to develop a model of trust in human-robot teams. It differs from previous modeling efforts in 3 different ways. First we investigated trust evolution in a human-robot collaborative task where we evaluated the human’s intervention in the robot’s actions and we sought to understand how human trust is affected if the robot asked for assistance. Using data from multiple participants across several trials, we estimate the parameters of an IOHMM model to explore how assistance-seeking, environmental complexity and prior experience influence human trust and action. The trust estimates with the estimated IOHMM parameters are compared with self-reported survey data. We then combine the IOHMM-based trust modeling with a POMDP-based decision-making framework to design an optimal assistance-seeking policy, comparing its effectiveness with a baseline trust-agnostic policy in further experiments. While IOHMMs and POMDPs have previously been used to model trust and design robot policy, to the best of our knowledge, they have not been applied in the context of assistance-seeking behavior within human-robot collaboration. Our formulation captures how a robot’s choices to seek or avoid help dynamically influence human trust over time.
@@
An early version of this work (DHM-ER-VS:22e, ) was presented at the 2023 American Control Conference. In addition to the results in (DHM-ER-VS:22e, ), this paper reports results from experiments with a significantly larger number of participants, reports uncertainty in the model fits, grounds the computational model by establishing an isomorphism between model predictions and survey responses, and provides a more detailed discussion.
@@ 2. Human-Supervised Robotic Object Collection Experiment
Figure 1. Experimental Setup. The top figure shows the world view of the environment available to the human supervisor. The bottom left and middle figures show the local view and end-effector view used by humans during teleoperation. The bottom right figure shows the operation status. The setup uses ROS-Gazebo (Gazebo, ) and resources available in (downs2022google, ; ROSMoMa, ).:
@@ 3.1. Background: IOHMM and POMDP
Consider a Markov chain with state $\{\bar{S}_{t}\in\bar{\mathcal{S}}\}_{t\in\mathbb{N}}$ with an input $\{u_{t}\in\mathcal{U}\}_{t\in\mathbb{N}}$, i.e., $\bar{S}_{t}$ is a random process such that the probability distribution of $\bar{S}_{t+1}$ is completely determined by $\bar{S}_{t}$ and $u_{t+1}$. Here, $\bar{\mathcal{S}}$ and $\mathcal{U}$ are some finite sets. In an IOHMM (536317, ), the state $\bar{S}_{t}$ is hidden and not directly observed, and only an output $\{y_{t}\in\mathcal{Y}\}_{t\in\mathbb{N}}$ is measured, which is a realization from an unknown probability distribution $\mathbb{P}(y_{t}|\bar{S}_{t},u_{t})$. Here, $\mathcal{Y}$ is some finite set. The objective of the IOHMM is to infer the hidden state $\bar{S}_{t}$ and learn its state transition matrix and $\mathbb{P}(\bar{S}_{t+1}|\bar{S}_{t},u_{t+1})$ from $y_{t}$ and $u_{t}$.
A POMDP (krishnamurthy2016partially, ) extends IOHMM by having some of the inputs controllable, while others are assumed Markovian and included in the state space. Consider the POMDP with augmented state $\{S_{t}\in\mathcal{S}\}_{t\in\mathbb{N}}$, controllable action sequence $\{a_{t}\in\mathcal{A}\}_{t\in\mathbb{N}}$, and observation sequence $\{o_{t}\in\mathcal{O}\}_{t\in\mathbb{N}}$. Along with the state set $\mathcal{S}$, action set $\mathcal{A}$, and observation set $\mathcal{O}$, the POMDP 7-tuple is completed by state transition probabilities $\mathbb{P}(S_{t+1}|S_{t},a_{t})$, observation probabilities $\mathbb{P}(o_{t}|S_{t},a_{t})$, a reward function $\mathcal{R}(S_{t},a_{t})$, and a discount factor $\gamma$. Note that state $S_{t}$ includes hidden state $\bar{S}_{t}$ and non-controlled inputs of the IOHMM. Additionally, $o_{t}$ includes perfect measurements of the non-controlled inputs and output $y_{t}$ of the IOHMM.
@@ 3.2. Trust-modulated Human Behavior Model
Building upon existing models in the literature (akash2017dynamic, ), we let the trust level $T_{t+1}$ at the end of trial $t$ as being influenced by the trust $T_{t}$ at the beginning of the trial, and the experience $E_{t+1}$ at the end of trial $t$ caused by the robot action $a_{t}$ at the complexity $C_{t}$, and the human action $o_{t}$, during trial $t$. Fig. 3 shows the IOHMM-based dynamic human behavior model, which includes the interconnection and influences of the hidden state, inputs, and observable human actions. Training the IOHMM using data sequences results in finding the probability values for trust dynamics $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ and human action selection $\mathbb{P}(o_{t}|T_{t},a_{t},C_{t})$
@@ 4. Estimated Human Behavioral Model
The collected data in each trial are human action $o$, robot action $a$, complexity $C$, and experience $E$. The data collected from all participants are combined and used to estimate the parameters of the human behavioral and trust dynamics model. Each time a shelf is cleared and objects are collected, human participants are asked to fill out a survey related to their trust. It should be noted that the survey responses are not used for learning the parameters of the human behavior model. The survey responses collected are discussed in Section 4.4. To learn the parameters of the IOHMM-based human behavior model, an extended version of the Baum-Welch algorithm is used (536317, ). [^3] Our code is available at https://github.com/donghaem14/Learning-algorithm-for-Input-Output-HMM-with-state-and-observation-inputs. The estimated parameters of the model are presented and discussed in the following section. For brevity and simplification of notation, the events are represented by their outcomes; for example, $E_{t}=E^{+}$ is shortened to $E^{+}_{t}$. Since the data was collected through in-person experiments, we were able to observe the behaviors of the participants and monitor them. Data collected from 6 participants are omitted due to behaviors observed during the experiment such as lack of attention, constantly checking their phone, and inability to follow the basic instructions of the experiment. This leaves 33 data sequences for model parameter estimation.

### v3

# When To Seek Help: Trust-Aware Assistance-Seeking in Human-Supervised Autonomy

@@ 1. Introduction
Autonomous systems are becoming increasingly prevalent and are being deployed and used in various sectors such as healthcare, agriculture, military, and transportation. Despite the increasing capabilities of autonomous systems, these systems still require human supervision and assistance when working in complex and uncertain environments(Selma2017; akash2020human; peterssupervisor2015; gupta2024structural; Boguslavskii2025). However, the increased capabilities of the robots are leading to novel landscapes of human-robot interaction (HRI) wherein the robots are deployed in roles requiring complex social behavior with human partners.
@@
Human trust in the robotic system is a key factor to be considered for effective collaboration between humans and robots in complex and uncertain environments. While trust has been traditionally studied in the context of collaborative human teams (mayer1995integrative), its influence in heterogeneous human-robot teaming is gaining prominence with the increased capabilities and deployment of robotic systems. Lee and See (lee2004trust) define trust in an autonomous agent as “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability”. Wagner and Arkin (wagner2011recognizing) generalize this definition to define trust as “belief held by the trustor that the trustee will act in a manner that mitigates the trustor’s risk in a situation in which the trustor has put its outcome at risk.” Trust in a robotic system is not only influenced by its physical design but also by the autonomous decisions and actions executed by the robot.
@@
While having low trust is undesirable as it leads to the disuse of a system, having high trust can also be undesirable as it can lead to misuse or abuse of the system (hoff2015trust; hancock2011meta; parasuraman1997humans). Therefore, for an effective design of human-robot teams, the calibration of the human partner’s trust in the robot’s abilities and capabilities is necessary. To this end, it is vital to understand the influence of various robot-related, environment-related, and human-related factors on the dynamic evolution of trust.
@@
The classical approach to studying the impact of robot reliability, transparency, and human workload on human trust treats trust as a static parameter (hoff2015trust). However, recent studies have focused on understanding the dynamic evolution of human-automation trust, focusing on factors such as the quality of robot performance, transparency of the robot’s operation, and the human agent’s attitude and characteristics, such as forgetfulness and self-confidence, on the trust dynamics (lee1992trust; lee1994trust; manzey2012human; desai2013impact; akash2017dynamic; yang2017evaluating; Wang2015DynamicRS; chen2020trust; akash2020human; williams2023computational). Hancock et al. (hancock2011meta) present a comprehensive review of various factors affecting human trust and categorize them into three major groups: robot-related, human-related, and environment-related factors. They discuss that human trust is most influenced by robot-related factors, which include robot reliability, performance, design, and abilities. This is followed by the environment-related factors, which include task complexity, type, and physical environment. Among the three main categories, human-related factors such as demographics, expertise, and personality have the least influence on trust in human-robot interaction.
@@
Deterministic linear models use a linear time-invariant dynamical system, where the states are the current trust level, cumulative trust level, and expectation bias (akash2017dynamic; yang2017evaluating; Wang2015DynamicRS). The input to these linear models includes human experience, which is determined by the performance and reliability of the robot. Variants of these models have been introduced to capture mutual trust between humans and robots (rahman2018mutual; Wang2014). These models have been used in several contexts, including exploring the impact of information transparency in reconnaissance missions (yang2017evaluating), investigating human reliance on driving assistance systems (akash2017dynamic), examining subtask allocation in collaborative assembly tasks (rahman2018mutual), and addressing scheduling challenges in teleoperation for underwater robotic navigation (Wang2014).
@@
Stochastic models of trust often take one of two forms. Stochastic linear models for the evolution of trust (azevedo2021real) rely on linear time-invariant dynamical systems with Gaussian measurement noise. In contrast, other stochastic models treat trust as a categorical latent variable and estimate its distribution based on observable factors like robot performance and human actions (OPTIMO; liao2024trust). An example of this class of models is the Input-Output Hidden Markov Model (IOHMM) with categorical states, which have been previously used to model human trust (akash2020human; DBLP:journals/corr/abs-2009-11890; williams2023computational). These models are particularly suitable for capturing interpretable internal states (e.g., “High Trust” or “Low Trust”). In addition, unlike linear Gaussian systems, which typically rely on richer or continuous signals such as self-reported trust on Likert scales (chen2020trust), gaze duration, or attention time (azevedo2021real), or the presence of a secondary task, the IOHMM provide flexibility for handling discrete observations such as human decisions (e.g., Rely or Intervene, Yes or No, Agree or Disagree) in both training and prediction. A comprehensive survey of different trust models is presented in (wang2023human). To optimize robot decision-making under uncertainty about internal human states, Partially Observable Markov Decision Process (POMDP) framework has been employed with trust as latent state influenced by robot actions, human responses, and contextual factors (chen2020trust; zahedi2023trust; akash2020human; 10.5555/2906831.2906852; DBLP:journals/corr/abs-2009-11890). It has been used in applications including scheduling in human-supervised robotics to reduce human interventions (chen2020trust) and monitoring load (zahedi2023trust), as well as generating optimal recommendations and explanations based on inferred human state (akash2020human; 10.5555/2906831.2906852). In this work, we combine IOHMM-based trust modeling with a POMDP-based decision-making framework.
@@
In addition to a dynamic trust model, real-time estimation of trust is vital for the design of trust-based policies that are robust to modeling uncertainties. Traditionally, human trust is assessed through self-reporting, such as surveys (schaefer2016measuring; lee1992trust; muir1989operators). However, a limitation of this approach is its impracticality for frequent trust survey completion during the experiment. Psycho-physiological measurements(gebru2022review) have also been used to infer the trust state; however, the sensing devices required to collect these measurements, e.g., an EEG head cap, limit their applicability. For a more detailed review of trust measurements, refer to (gebru2022review). These challenges have prompted researchers to rely purely on context and behavioral data, such as human actions, to estimate trust. This approach has been implemented and tested in several contexts such as reconnaissance mission (akash2020human), autonomous driving (azevedo2021real; DBLP:journals/corr/abs-2009-11890), and supervision(OPTIMO). We focus on similar methodology for estimating human trust: subjective measures and behavioral data.
@@
In this paper, in the context of human-supervised object collection tasks, we design and validate trust-aware policies for an autonomous system seeking assistance from a human supervisor. In contrast to existing works, our focus is on a scenario where the robot can seek assistance from the human supervisor, preventing a loss of trust and possibly repairing it. The major contributions of this work are the following. First, we develop a novel experiment to explore how a robot’s requests for assistance influence human trust. Second, we use an Input-Output Hidden Markov Model (IOHMM) (536317) to capture the evolution of the hidden human trust state. In our formulation, the inputs are the robot’s action, environmental complexity, and the outcome or experience in the previous trial, characterized by the success or failure of the robot in accomplishing its goals, and the output is the human action. The present research sought to develop a model of trust in human-robot teams. It differs from previous modeling efforts in 3 different ways. First we investigated trust evolution in a human-robot collaborative task where we evaluated the human’s intervention in the robot’s actions and we sought to understand how human trust is affected if the robot asked for assistance. Using data from multiple participants across several trials, we estimate the parameters of an IOHMM model to explore how assistance-seeking, environmental complexity and prior experience influence human trust and action. The trust estimates with the estimated IOHMM parameters are compared with self-reported survey data. We then combine the IOHMM-based trust modeling with a POMDP-based decision-making framework to design an optimal assistance-seeking policy, comparing its effectiveness with a baseline trust-agnostic policy in further experiments. While IOHMMs and POMDPs have previously been used to model trust and design robot policy, to the best of our knowledge, they have not been applied in the context of assistance-seeking behavior within human-robot collaboration. Our formulation captures how a robot’s choices to seek or avoid help dynamically influence human trust over time.
@@
An early version of this work (DHM-ER-VS:22e) was presented at the 2023 American Control Conference. In addition to the results in (DHM-ER-VS:22e), this paper reports results from experiments with a significantly larger number of participants, reports uncertainty in the model fits, grounds the computational model by establishing an isomorphism between model predictions and survey responses, and provides a more detailed discussion.
@@ 2. Human-Supervised Robotic Object Collection Experiment
Figure 1. Experimental Setup. The top figure shows the world view of the environment available to the human supervisor. The bottom left and middle figures show the local view and end-effector view used by humans during teleoperation. The bottom right figure shows the operation status. The setup uses ROS-Gazebo (Gazebo) and resources available in (downs2022google; ROSMoMa).:
@@ 3.1. Background: IOHMM and POMDP
Consider a Markov chain with state $\{\bar{S}_{t}\in\bar{\mathcal{S}}\}_{t\in\mathbb{N}}$ with an input $\{u_{t}\in\mathcal{U}\}_{t\in\mathbb{N}}$, i.e., $\bar{S}_{t}$ is a random process such that the probability distribution of $\bar{S}_{t+1}$ is completely determined by $\bar{S}_{t}$ and $u_{t+1}$. Here, $\bar{\mathcal{S}}$ and $\mathcal{U}$ are some finite sets. In an IOHMM (536317), the state $\bar{S}_{t}$ is hidden and not directly observed, and only an output $\{y_{t}\in\mathcal{Y}\}_{t\in\mathbb{N}}$ is measured, which is a realization from an unknown probability distribution $\mathbb{P}(y_{t}|\bar{S}_{t},u_{t})$. Here, $\mathcal{Y}$ is some finite set. The objective of the IOHMM is to infer the hidden state $\bar{S}_{t}$ and learn its state transition matrix and $\mathbb{P}(\bar{S}_{t+1}|\bar{S}_{t},u_{t+1})$ from $y_{t}$ and $u_{t}$.
A POMDP (krishnamurthy2016partially) extends IOHMM by having some of the inputs controllable, while others are assumed Markovian and included in the state space. Consider the POMDP with augmented state $\{S_{t}\in\mathcal{S}\}_{t\in\mathbb{N}}$, controllable action sequence $\{a_{t}\in\mathcal{A}\}_{t\in\mathbb{N}}$, and observation sequence $\{o_{t}\in\mathcal{O}\}_{t\in\mathbb{N}}$. Along with the state set $\mathcal{S}$, action set $\mathcal{A}$, and observation set $\mathcal{O}$, the POMDP 7-tuple is completed by state transition probabilities $\mathbb{P}(S_{t+1}|S_{t},a_{t})$, observation probabilities $\mathbb{P}(o_{t}|S_{t},a_{t})$, a reward function $\mathcal{R}(S_{t},a_{t})$, and a discount factor $\gamma$. Note that state $S_{t}$ includes hidden state $\bar{S}_{t}$ and non-controlled inputs of the IOHMM. Additionally, $o_{t}$ includes perfect measurements of the non-controlled inputs and output $y_{t}$ of the IOHMM.
@@ 3.2. Trust-modulated Human Behavior Model
Building upon existing models in the literature (akash2017dynamic), we let the trust level $T_{t+1}$ at the end of trial $t$ as being influenced by the trust $T_{t}$ at the beginning of the trial, and the experience $E_{t+1}$ at the end of trial $t$ caused by the robot action $a_{t}$ at the complexity $C_{t}$, and the human action $o_{t}$, during trial $t$. Fig. 3 shows the IOHMM-based dynamic human behavior model, which includes the interconnection and influences of the hidden state, inputs, and observable human actions. Training the IOHMM using data sequences results in finding the probability values for trust dynamics $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ and human action selection $\mathbb{P}(o_{t}|T_{t},a_{t},C_{t})$
@@ 4. Estimated Human Behavioral Model
The collected data in each trial are human action $o$, robot action $a$, complexity $C$, and experience $E$. The data collected from all participants are combined and used to estimate the parameters of the human behavioral and trust dynamics model. Each time a shelf is cleared and objects are collected, human participants are asked to fill out a survey related to their trust. It should be noted that the survey responses are not used for learning the parameters of the human behavior model. The survey responses collected are discussed in Section 4.4. To learn the parameters of the IOHMM-based human behavior model, an extended version of the Baum-Welch algorithm is used (536317). [^3] Our code is available at https://github.com/donghaem14/Learning-algorithm-for-Input-Output-HMM-with-state-and-observation-inputs. The estimated parameters of the model are presented and discussed in the following section. For brevity and simplification of notation, the events are represented by their outcomes; for example, $E_{t}=E^{+}$ is shortened to $E^{+}_{t}$. Since the data was collected through in-person experiments, we were able to observe the behaviors of the participants and monitor them. Data collected from 6 participants are omitted due to behaviors observed during the experiment such as lack of attention, constantly checking their phone, and inability to follow the basic instructions of the experiment. This leaves 33 data sequences for model parameter estimation.